OpenStudy (rebecca1588):

Is each situation a linear or a nonlinear model? 1. The stopping distance for a car is given by the equation D=v2/15.7 where D is the distance a car travels once the brakes are applied and v is the car’s velocity. A. linear model B. nonlinear model

OpenStudy (anonymous):

Let x stand for time, y stand for velocity. We are given the points (2,50), (6, 54). We can make a line using the slope intercept form y = mx + b. slope is (54 - 50)/(6-2) = 4/4 = 1 y = 1x + b plug in point (2,50) to find b 50 = 1(2) + b 50-2 = b 48 = b the equation is y = 1x + 48 Make standard form. x - y = -48

OpenStudy (anonymous):

Hi

OpenStudy (rebecca1588):

@superdavesuper Could you help me

OpenStudy (wwb00):

Sorry, I thought the v term was squared.

Directrix (directrix):

If this is D=v^2/15.7, then it is not a linear model because linear models do not have a power of x greater than one.

OpenStudy (rebecca1588):

Thanks :)

Directrix (directrix):

You are welcome.